# Boundary Exploration of Next Best View Policy in 3D Robotic Scanning

### v1

@@ Abstract
The Next Best View (NBV) problem is a pivotal challenge in 3D robotic scanning, with the potential to greatly improve the efficiency of object capture and reconstruction. Current methods for determining the NBV often overlook view overlaps, assume a virtual origin point for the camera’s focus, and rely on voxel representations of 3D data. To address these issues and improve the practicality of scanning unknown objects, we propose an NBV policy in which the next view explores the boundary of the scanned point cloud, and the overlap is intrinsically considered. The scanning distance or camera working distance is adjustable and flexible. To this end, a model-based approach is proposed where the next sensor positions are searched iteratively based on a reference model. A score is calculated by considering the overlaps between newly scanned and existing data, as well as the final convergence. Additionally, following the boundary exploration idea, a deep learning network, Boundary Exploration NBV network (BENBV-Net), is designed and proposed, which can be used to predict the NBV directly from the scanned data without requiring the reference model. It predicts the scores for given boundaries, and the boundary with the highest score is selected as the target point of the next best view. BENBV-Net improves the speed of NBV generation while maintaining the performance of the model-based approach. Our proposed methods are evaluated and compared with existing approaches on the ShapeNet, ModelNet, and 3D Repository datasets. Experimental results demonstrate that our approach outperforms others in terms of scanning efficiency and overlap, both of which are crucial for practical 3D scanning applications. The related code is released at github.com/leihui6/BENBV.
@@ I Introduction
Figure 1: Overview of our proposed approach: Starting with an initial scan, the boundary of the raw point cloud is detected and clustered to generate proposed views. In the model-based approach, the NBV is searched and selected based on the reference model. In the model-free approach, we predict scores for given boundaries and select the one with the highest score to be the NBV.
@@
The NBV problem has been widely explored using 2D images, but 2D approaches are often limited by the lack of depth information and ambiguity in spatial relationships. In contrast, depth data or point clouds inherently capture rich and geometric details about the target and scene, making them more effective for NBV planning in complex environments. NBV methods utilizing 3D data from point clouds can broadly be categorized into model-based and model-free approaches [5]. Model-based methods leverage pre-built or provided reference models of target objects to predict the next best view, with the reference model involved in each step. In contrast, model-free approaches generate the NBV without prior knowledge of the object’s reference model, relying solely on the information acquired during the scanning process. Most existing NBV methods utilize mesh or voxel representations of the retrieved 3D point cloud, which are well-suited for large outdoor scenes with low precision requirements. However, these methods require additional preprocessing, such as surface reconstruction and voxelization. These processes increases computational and storage costs while offering lower precision for indoor environments or small objects.
On the other hand, existing methods typically rely on empirically designed view spaces, such as a hemisphere around the target object. However, these approaches limit their generalizability to novel objects. Additionally, they overlook the optimization of the sensor’s working distance, assuming the origin point is at the center of the hemisphere, which can lead to variations in imaging quality. Furthermore, while many existing methods assess scanning efficiency based on coverage, they often overlook the significance of overlap between newly acquired data and existing datasets. This overlap is crucial for properly aligning new data with the current dataset, especially in cases where camera movement deviates from the expected position and orientation.
To address these challenges and improve NBV selection performance, we introduce a boundary exploration NBV policy. This approach identifies boundary points from the scanned data as candidate views, where camera positions are defined, and the camera’s focus is directed toward the boundaries. First, a model-based method is proposed, which iteratively searches for and selects the NBV with the highest score. This score balances overlap and coverage to optimize scanning efficiency. Second, we present a model-free method utilizing a deep learning framework, the Boundary Exploration NBV Network (BENBV-Net). BENBV-Net predicts scores for the boundary points of the point cloud and selects the point with the highest score as the NBV. This approach eliminates the need for a reference model while improving the inference speed for NBV selection.
@@
    We propose an intuitive NBV policy that explores the boundaries of point clouds directly using the raw data as input. This approach accounts for both the overlap between newly captured and existing data and the overall coverage. Additionally, it allows for an adjustable camera working distance, making it adaptable to various 3D sensors.
@@
    We propose a model-based approach, where a reference model is used to search and select the view with the highest score as the NBV. Additionally, we introduce a learning-based framework, BENBV-Net, which does not require a reference model. Instead, it takes the scanned point cloud and boundaries as input and predicts scores for each boundary.
@@
    The efficiency and effectiveness of our approach are thoroughly evaluated against traditional methods on datasets such as ShapeNet, ModelNet, and 3D Repository. Furthermore, the proposed BENBV-Net showcases generalization capabilities on unseen test data and novel objects, achieving performance comparable to the model-based approach while significantly reducing computational time.
@@ II Related Work
To this end, we propose an intuitive NBV policy that explores the boundary of the acquired point cloud. The overlap is naturally accounted for, as the next view can always be oriented towards the existing data. Most importantly, our method allows for adjusting the distance between the target surface and the 3D acquisition device to optimize scanning, rather than assuming the object’s centroid or original position as the camera’s focal point.
@@ III-A Problem Definition
The $F$ here is defined by the coverage ratio $\text{C}=\mathopen{|}s_{i}\mathclose{|}/\mathopen{|}S\mathclose{|}$ and the overlap ratio $\text{O}=\mathopen{|}p_{o}\mathclose{|}/\mathopen{|}s_{i}\mathclose{|}$, where $S$ represents the number of points in the reference model, $S_{i}$ is the number of points in the retrieved point cloud, and $\mathopen{|}p_{o}\mathclose{|}$ represents number of overlapping points between the current scanned data and the existing data. In a simulated environment, $S$ is known and predefined, and $s_{i}$ is generated within a physics-based simulation engine. However, for real-world applications, $s_{i}$ is derived from actual 3D sensor data, and $S$ is typically unknown in advance due to the novelty of the target object.
@@ III-B Framework Overview
Given the captured point cloud, the general idea in this paper is to explore the unseen regions by following its boundary. However, not all boundaries contribute meaningfully to high coverage and overlap. Therefore, we first introduce a model-based approach that requires a reference model, where the NBV is determined through a search process, and the view with the highest score is selected as the next best view. We then present a model-free approach, where we develop and propose a deep learning network that learns the latent space mapping between the designated boundary and the point cloud. This enables the network to predict the most optimal NBV in a time-efficient manner. The overall framework is illustrated in Figure 1.
@@ III-C Boundary exploration for NBV policy
The boundaries of retrieved 3D representation data have been examined in previous studies [22, 12], which utilize either triangle surfaces or density-based methods to explore edges. In our study, we compute the boundary directly from the raw point cloud data using the Angle Criterion method proposed by [23] where the angle threshold is set 120 degree. Let $B_{i}$ represent the boundary points for the $i$-th scanned data. To reduce the large number of boundary points and simplify the process, we use the K-Means algorithm to cluster the boundary points into 20 clusters. From each cluster, we randomly select one point as $V^{tar}_{i}$.
The normals for the scanned data are estimated as $s_{i}^{N\times 6}$and are used to determine the direction of the camera position. Specifically, the normals are oriented such that their direction aligns toward the camera position, ensuring the angle between the normals and the camera direction is less than 90. For each $V^{tar}_{i}$, we establish a local orthogonal coordinate system (u-v-n). In this system, n represents the normal vector at $V^{tar}_{i}$, while u is defined by
@@
where $\theta$ is randomly set to $-45^{\circ},0^{\circ}\text{ or }45^{\circ}$. Given the scanning distance $d$ which is adjustable and can be changed during scanning process, the camera position is determined as
@@
Figure 2: u-v-n: the $V^{tar}_{i}$ is marked in black and the $\bar{V}^{tar}_{i}$ is denoted in red. The camera position is defined on the right, with $d$ indicating the specified distance.
Given the proposed $V^{cam}_{i}$ and $V^{tar}_{i}$, we need to assign a score to each potential view by considering both coverage and overlap, and the view with the highest $s$ is selected as the next best view. The score for each view is calculated as
@@
where $O_{i}$ and $C_{i}$ represent the overlap and coverage values at the $i$-th view. The weight $W_{c}$ is defined as
$W_{c}=\frac{1}{1+e^{-10\cdot(C_{i}-0.6)}}$ (6)
Coverage becomes increasingly important, while overlap becomes less significant as the scanning process progresses. Specifically, overlap is weighted more heavily than coverage until the current coverage reaches 60%.
The workflow for searching the NBV is detailed in Algorithm. 1. For the initial view, $V^{cam}_{best}$ is chosen from the surface of a sphere with radius $d$, while $V^{tar}_{best}$ is set as the origin point $(0,0,0)$. The simulation engine used in simulator can provide the virtual scanned data. The detailed searching process is outlined in Algorithm. 2. The dataset in Algorithm 1 is constructed for the deep learning network we developed, where the proposed views, their corresponding point clouds, and the associated scores are stored.
@@
10:     ${\rm existedData}\leftarrow{\rm existedData+scannedData}$
11:     ${\rm(V^{cam}_{best},\ V^{tar}_{best}),viewList,SList}\leftarrow{\rm searchNBV}(\rm...)$
12:     ${\rm Dataset}\leftarrow{\rm(existedData,viewList,SList)}$
@@
9:     ${\rm S}\leftarrow\left(1-W_{c}\right)\cdot O+W_{c}\cdot C$
@@ III-D Deep Learning Architecture
[…]
We propose BENBV-Net, a deep neural network for hierarchical feature learning and NBV score prediction from 3D point clouds. The network employs two parallel encoders: a point feature encoder for spatial coordinates and a normal feature encoder for surface normals, both utilizing convolutional layers with batch normalization and ReLU activation. The P feature follows a structure similar to PointNet, processing point cloud data in a permutation-invariant manner.
Boundary features are extracted via fully connected layers, while contextual features, including point density and view order, are fused using a learnable density and view order fusion module. These features are combined with global representations through residual blocks for multi-scale refinement. A multi-head self-attention mechanism captures long-range dependencies.
The prediction head computes per-point NBV scores using fully connected layers with dropout regularization. The entire network is trained end-to-end from raw point cloud data. An overview is shown in Figure 3.
Figure 3: Our developed deep learning framework, BENBV-Net, for the NBV policy takes the raw point cloud, boundary points, and the corresponding context as input. The network first extracts the global feature of the current point cloud, which is then fused with the global feature from the boundary points, along with the features extracted from the context. These combined features are used to predict a score for each boundary point, with the highest-scoring boundary selected as the NBV.
@@ IV-A Experiment Setup
We use PyBullet [24] as the simulation platform, with the camera configuration set to a near distance of 0.01 m, a far distance of 5 m, and a captured image size of $1280\times 720$ pixels, with a $70^{\circ}$ field of view (FOV). The scanning distance for 3D camera is set to 1.2 m. All experiments were conducted on a system running Ubuntu 24.04, equipped with an Intel i9 processor and an NVIDIA RTX 3090 GPU.
@@ IV-B Dataset
We evaluate and train our model-based method and BENBV-Net using the ShapeNet [25], ModelNet40 [26] datasets, and 3D Repository from Stanford University 111graphics.stanford.edu/data/3Dscanrep/ and Georgia Tech 222sites.cc.gatech.edu/projects/large_models/, with an example shown in Figure 4. The ShapeNet and ModelNet40 datasets are used for both training and evaluation, while the 3D Repository is reserved exclusively for testing. Specifically, we use ShapeNetV1, selecting 20 models per category for training and 15 models per category for testing. Similarly, for ModelNet40, which includes 40 categories of CAD-generated meshes, we select 20 models per category for training and 15 for testing. For each model, 16 scans are extracted as trainable data. This setup results in a total of over 24,000 training samples from ShapeNetV1 and ModelNet40, 18,000 testing samples, and 128 testing samples from the 3D Repository. The datasets we built and utilized are available at huggingface.co/datasets/Leihui/NBV. The model used in the simulation environment for scanning is sampled as points using Poisson Disk Sampling [27].
Figure 4: An example from the 3D Repository dataset, which consists of classic 3D data widely used in the 3D research field.
@@ IV-D Evaluation Metrics
The performance of the comparison methods is evaluated using the following metrics: 1) final coverage, 2) overlap between the current view and existing data, 3) Chamfer Distance (CD), 4) Hausdorff Distance (HD), and 5) scanning efficiency. Specifically, we calculate the average of the initial overlap values, as early overlaps are more critical and challenging to optimize in practical scenarios. Furthermore, we evaluate scanning quality by comparing the final retrieved points with the original model using Chamfer and Hausdorff distances. The Chamfer Distance measures the average distance between two point sets, providing an overall indication of their similarity, while the Hausdorff Distance captures the maximum distance from any point in one set to its nearest neighbor in the other, emphasizing worst-case alignment errors. Additionally, scanning efficiency, which aligns with the objectives of the NBV policy, is evaluated by
@@
where $\rm v$ represent the number of views required to achieve 90% coverage, and $\rm c\in[0,1]$ denote the final coverage ratio. A higher $e$ value indicates greater efficiency in the NBV policy. Specifically, it reflects the ability of the policy to achieve a high coverage with fewer views.

### v2

# Boundary Exploration for Next Best View Policy in 3D Robotic Scanning

@@ Abstract
The Next Best View (NBV) problem is a pivotal challenge in 3D robotic scanning, with the potential to significantly improve the efficiency of object capture and reconstruction. Existing methods for determining the NBV often overlook view overlap, assume a fixed virtual origin for the camera, and rely on voxel-based representations of 3D data. To address these limitations and enhance the practicality of scanning unknown objects, we propose an NBV policy in which the next view explores the boundary of the scanned point cloud, with overlap intrinsically considered. The scanning or working distance of the camera is user-defined and remains flexible throughout the process. To this end, we first introduce a model-based approach in which candidate views are iteratively proposed based on a reference model. Scores are computed using a carefully designed strategy that accounts for both view overlap and convergence. In addition, we propose a learning-based method, the Boundary Exploration NBV Network (BENBV-Net), which predicts the NBV directly from the scanned data without requiring a reference model. BENBV-Net estimates scores for candidate boundaries, selecting the one with the highest score as the target for the next best view. It offers a significant improvement in NBV generation speed while maintaining the performance level of the model-based approach. We evaluate both methods on the ShapeNet, ModelNet, and 3D Repository datasets. Experimental results demonstrate that our approach outperforms existing methods in terms of scanning efficiency, final coverage, and overlap stability, all of which are critical for practical 3D scanning applications. The related code is available at github.com/leihui6/BENBV.
@@ I Introduction
Figure 1: Overview of the proposed approach. Starting from an initial scan, the boundary of the raw point cloud is detected and clustered to generate candidate views. In the model-based approach, the NBV is selected through a search on the reference model. In the model-free approach, scores are predicted for each boundary, and the highest-scoring view is chosen as the NBV.
@@
The NBV problem has been widely explored using 2D images, but 2D approaches are often limited by the lack of depth information and ambiguity in spatial relationships. In contrast, depth data or point clouds inherently capture rich and geometric details about the target and scene, making them more effective for NBV planning in complex environments. NBV methods utilizing 3D data from point clouds can broadly be categorized into model-based and model-free approaches [5]. Model-based methods leverage pre-built or provided reference models of target objects to predict the next best view, with the reference model involved in each step. In contrast, model-free approaches generate the NBV without prior knowledge of the object’s reference model, relying solely on the information acquired during the scanning process. In addition, most existing NBV methods utilize mesh or voxel representations of the retrieved 3D point cloud, which are well-suited for large outdoor scenes with low precision requirements. However, these methods require additional preprocessing, such as surface reconstruction and voxelization. These processes increase computational and storage costs while offering lower precision for indoor environments or small objects.
On the other hand, existing methods typically rely on empirically designed view spaces, such as a hemisphere around the target object. However, these approaches limit their generalizability to novel objects. Additionally, they overlook the optimization of the sensor’s working distance, assuming the origin point is at the center of the hemisphere, which can lead to variations in imaging quality. Furthermore, while many existing methods assess scanning efficiency based on coverage, they often overlook the significance of overlap between newly acquired data and existing datasets. This overlap is crucial for properly aligning new data with the existing dataset, especially during the initial scanning stage or when the camera deviates from its expected position and orientation.
To address these challenges and improve NBV selection performance, we introduce a boundary exploration NBV policy. This approach identifies candidate views from the boundary points of the raw accumulated point cloud, directing the camera’s focus toward these boundaries. Specifically, we first propose a model-based method that iteratively searches for and selects the NBV with the highest score. This score balances overlap and coverage to optimize scanning efficiency. Second, we present a model-free method based on a deep learning framework, the Boundary Exploration NBV Network (BENBV-Net). BENBV-Net predicts scores for the boundary points of the point cloud, and the point with the highest score is selected as the NBV. This model eliminates the need for a reference model while significantly improving the inference speed of NBV selection.
@@
    We propose an intuitive NBV policy that directly explores the boundaries of point clouds using raw input data. The approach considers both view overlap and coverage, ensuring stable overlap and high efficiency during data acquisition.
@@
    We introduce a model-based approach in which a reference model is used to iteratively select the view with the highest score as the NBV. In addition, we present a learning-based framework, BENBV-Net, which eliminates the need for a reference model by taking the scanned point cloud and its boundaries as input.
@@
    The efficiency and effectiveness of our methods are evaluated against traditional approaches on public datasets. BENBV-Net demonstrates strong generalization to unseen data and novel objects, achieving performance comparable to the model-based method while significantly reducing computational time.
@@ II Related Work
The proposed NBV policy extends the accumulated data by exploring the boundary of the point cloud. Overlap and coverage are jointly considered through a thoughtfully designed strategy. Additionally, our method allows flexible adjustment of the distance between the target surface and the 3D acquisition device, optimizing the scanning process without assuming the object’s centroid or a fixed origin as the camera’s focal point.
@@ III-A Problem Definition
The metric $F$ is defined based on coverage and overlap ratios. Specifically, the coverage is calculated as $\text{C}=\lvert s_{i}\rvert/\lvert S\rvert$, where $S$ denotes the total number of points in the reference model, and $s_{i}$ represents the number of points in the retrieved point cloud. The overlap ratio is given by $\text{O}=\lvert p_{o}\rvert/\lvert s_{i}\rvert$, where $p_{o}$ is the number of overlapping points between the newly scanned data and the accumulated point cloud.
In a simulated environment, the target object is predefined, allowing $S$ to be known, and $s_{i}$ can be generated using a physics-based simulation engine. In contrast, for real-world applications, $s_{i}$ is obtained from actual 3D sensor data, while $S$ is typically unknown due to the novelty of the target object.
@@ III-B Framework Overview
Given the captured point cloud, the general idea of this work is to explore unseen regions by following its boundary. However, not all boundary regions contribute meaningfully to achieving high coverage and overlap, and are therefore not always worth exploring or extending. To address this, we first introduce a model-based approach that utilizes a reference model. In this approach, the next-best view (NBV) is determined through a search process, and the view with the highest score is selected as the optimal NBV.
We then propose a model-free approach, in which a deep learning network is trained to predict the view selection score without relying on a reference model. The network is trained on a dataset generated by the model-based approach and learns a latent space mapping between the designated boundary and the corresponding point cloud. The overall framework is illustrated in Figure. 1.
@@ III-C Boundary exploration for NBV policy
The boundaries of retrieved 3D representation data have been examined in previous studies [22, 12], which utilize either triangle surfaces or density-based methods to explore edges. In our study, we compute the boundary directly from the raw point cloud data using the Angle Criterion method proposed by [23] where the angle threshold is set 120 degree. Let $B_{i}$ represent the boundary points for the $i$-th scanned data. To reduce the large number of boundary points and simplify the process, we use the K-Means algorithm to cluster the boundary points into 20 clusters. Within each cluster, the point closest to the cluster centroid is selected and denoted as $V^{\text{tar}}_{i}$. In addition, the set of $V^{\text{tar}}_{i}$ is resorted using a greedy nearest neighbor traversal, where each subsequent point is selected as the closest unvisited one to the current point.
The normals for the scanned data are estimated as $s_{i}^{N\times 6}$and are involved to determine the direction of the camera position. More specifically, a local orthogonal coordinate system (u-v-n) is established at $V^{\text{tar}}_{i}$, where n denotes its normal vector, and u is defined by
@@
where $\theta$ is set to $-45^{\circ}$, $0^{\circ}$, or $45^{\circ}$ in sequence. An angle of $0^{\circ}$ is used to scan the boundary vertically, while $-45^{\circ}$ and $45^{\circ}$ are chosen to address cases where the boundary lies near the edge, which may otherwise result in empty data. Given the scanning distance $d$, which is user-defined and adjustable for each scan, the camera position is determined as follow
@@
Figure 2: u-v-n: the $V^{tar}_{i}$ is marked in black and the $\bar{V}^{tar}_{i}$ is denoted in red. The camera position is defined on the right where $d$ is the specified distance.
Given the proposed $V^{\text{cam}}_{i}$, $V^{\text{tar}}_{i}$, and the simulated captured data, a score is assigned to each view by evaluating both coverage and overlap. The view with the highest score is selected as the next-best view. The score at the $i$-th view is calculated as
@@
where $O_{i}$ and $C_{i}$ represent the estimated overlap and coverage values at the $i$-th view. The weight $W_{c}$ is defined as
$W_{c}=\frac{1}{1+e^{-10\cdot(C_{i}-0.5)}}$ (6)
where $C_{i}$ denotes the coverage of the current view. It is assumed that coverage becomes increasingly important, while overlap becomes less significant as the scanning process progresses. Specifically, overlap is weighted more heavily than coverage until the current coverage exceeds 50%.
Furthermore, since the overlap values fluctuate across different views, a response function $OR(\cdot)$ is designed to smoothly map the overlap values to continuous values as
[…]
where the overlap range between 0.4 and 0.5 is considered the most representative, while values that are too low ($<$ 0.4) or too high ($>$ 0.5) are penalized. The response is zero at both extremes, specifically when $O_{i}=0$ and $O_{i}=1$.
The workflow for searching the NBV is detailed in Algorithm. 1. For the initial view, $V^{cam}_{best}$ is chosen from the surface of a sphere with radius $d$, while $V^{tar}_{best}$ is set as the origin point $(0,0,0)$. The scanned data is provided by the simulation engine. The detailed searching process is outlined in Algorithm. 2. The dataset generated in Algorithm. 1 is constructed for the deep learning network we developed, where the proposed views (viewList), corresponding point clouds (existedData), and the associated scores (SList) are collected.
@@
10:     ${\rm(V^{cam}_{best},\ V^{tar}_{best}),viewList,SList}\leftarrow{\rm searchNBV}(\rm...)$
11:     ${\rm Dataset}\leftarrow{\rm(existedData,viewList,SList)}$
12:     ${\rm existedData}\leftarrow{\rm existedData+scannedData}$
@@
9:     ${\rm S}\leftarrow W_{c}^{i}\cdot C+\left(1-W_{c}^{i}\right)\cdot OR(O)$
@@ III-D Deep Learning Architecture
The goal of the learning-based approach is to predict the NBV based on the detected boundaries, particularly in 3D automatic scanning tasks involving novel objects for which reference models are unavailable. The training dataset includes the point cloud (P), the detected boundaries (B), and their corresponding scores, which are used as prediction targets. To enable the network to efficiently learn to select the best view, the training supervision pair for the BENBV-Net is defined as (P, B, C) where C consists of the density for B and view order.
We propose BENBV-Net, a deep neural network for hierarchical feature learning and NBV score prediction from 3D point clouds. The network employs two parallel encoders: a point feature encoder for spatial coordinates and a normal feature encoder for surface normals, both utilizing convolutional layers with batch normalization and ReLU activation. The P feature follows a structure similar to PointNet, processing point cloud data in a permutation-invariant manner. Boundary features are extracted via fully connected layers, while contextual features, including point density and view order, are fused using a learnable density and view order fusion module. These features are combined with global representations through residual blocks for multi-scale refinement. A multi-head self-attention mechanism captures long-range dependencies.
The prediction head computes per-point NBV scores using fully connected layers with dropout regularization. The entire network is trained end-to-end from raw point cloud data. An overview is shown in Figure. 3.
Figure 3: Our developed deep learning framework: BENBV-Net takes the raw point cloud, boundary points, and the corresponding context as input. The network first extracts the global feature of the current point cloud, which is then fused with the global feature from the boundary points, along with the features extracted from the context. These combined features are used to predict scores for each boundary point, and the highest-scoring boundary is selected as the NBV.
@@ IV-A Experiment Setup
We use PyBullet [24] as the simulation platform, with the camera configuration set to a near distance of 0.01 m, a far distance of 5 m, and a captured image size of $1280\times 720$ pixels, with a $70^{\circ}$ field of view (FOV). The scanning distance for the 3D camera is set to 1.2 meters, which is an optimal distance for most low-cost cameras. The training of BENBV-Net and all experiments were conducted on a system running Ubuntu 24.04, equipped with an Intel i9 processor and an NVIDIA RTX 3090 GPU.
@@ IV-B Dataset
We evaluate and train our model-based method and BENBV-Net using the ShapeNet [25], ModelNet40 [26] datasets, and 3D Repository from Stanford University 111graphics.stanford.edu/data/3Dscanrep/ and Georgia Tech 222sites.cc.gatech.edu/projects/large_models, with an example shown in Figure 4. The ShapeNet and ModelNet40 datasets are used for both training and evaluation, while the 3D Repository is reserved exclusively for evaluation. Specifically, we use ShapeNetV1, selecting 20 models per category for training and 15 models per category for testing. Similarly, for ModelNet40, which includes 40 categories of CAD-generated meshes, we select 20 models per category for training and 15 for testing. For each model, 16 scans are extracted as trainable data. This setup results in a total of over 24,000 training samples from ShapeNetV1 and ModelNet40, 18,000 testing samples, and 128 testing samples from the 3D Repository. The datasets we built and utilized are available online333huggingface.co/datasets/Leihui/NBV. The scanning model used in our simulation platform is sampled as points using Poisson disk sampling [27].
Figure 4: An example from the 3D Repository dataset, which contains classic 3D models widely used in the research community. Additionally, point clouds from ShapeNet and ModelNet40 are also included.
@@ IV-D Evaluation Metrics
The performance of the comparison methods is evaluated using the following metrics: 1) final coverage, 2) stability of overlap between the newly current view and existing data, 3) Chamfer Distance (CD), 4) Hausdorff Distance (HD), and 5) scanning efficiency. Specifically, we calculate the average of the initial overlap values, as early overlaps are more critical and challenging to optimize in practical scenarios. The coefficient of variation (CV) of the overlap is reported for the first three scans. Furthermore, the scanning quality is evaluated by comparing the final retrieved points with the original model using Chamfer and Hausdorff distances. The Chamfer Distance measures the average distance between two point sets, providing an overall indication of their similarity, while the Hausdorff Distance captures the maximum distance from any point in one set to its nearest neighbor in the other, emphasizing worst-case alignment errors. Additionally, scanning efficiency, which aligns with the objectives of the general next best view policy, is evaluated by
@@
where $\rm v$ represents the number of views required to first reach or exceed 90% coverage, and $\rm c\in[0,1]$ denotes the final coverage ratio. A higher $e$ value indicates greater efficiency of the NBV policy, meaning that high coverage is achieved with fewer views.